# MobileVLM: A Fast, Strong and Open Vision Language Assistant for Mobile Devices

### v1

@@ 2.3 VLMs
| Model | Vision Encoder | Language Model | Cross-modality Design | Multimodal Training Corpora |
| CLIP [97] | ViT, ResNet | Transformer | Linear-Projection | WebImageText [97] (400M image-text pairs) |
| BLIP [67] | ViT | MED∗ | Cross-Attention | COCO [73], VG [62], CC3M [103], CC12M [13], LAION [102] |
| BLIP-2 [66] | CLIP/Eva-CLIP ViT | OPT, Flan-T5 | Q-Former | same as BLIP |
| InstructBLIP [30] | ViT-G/14 | Flan-T5, Vicuna | Q-Former w/ FC | 13 held-in out of 26 public datasets |
| Flamingo [2] | NFNet-F6 [59] | Chinchilla | Perceiver-Resampler [57] | M3W(43M), ALIGN(1.4B) [58], LTIP (312M), VTP (27M) |
| LLaVA [76] | CLIP ViT-L/14 | Vicuna 7B/13B | Linear-Pprojection | filtered CC-595K from CC3M [103], LLaVA-Instruct-158K |
| LLaVA-1.5 [74] | CLIP ViT-L@336 | Vicuna-7B/13B | MLP | a subset of InstructBLIP (1.2M) |
| MiniGPT-4 [133] | EVA-CLIP ViT-G/14 | Vicuna-7B | Q-Former | LAION, CC, SBU [91] |
| Shikra [15] | CLIP ViT-L/14 | Vicuna-7/13B | FC-layer | Flickr30K [95], RefCOCO [61], VG, Visual-7W [85] |
| mPLUG-Owl [126] | CLIP ViT-L/14 | LLaMA-7B | Cross-Attention | LAION-400M, COYO-700M [11], CC, COCO |
| KOSMOS-2 [94] | VLMo [6] | MAGNETO [119] | Perceiver-Resampler | GRIT (curated with LAION-2B, COYO-700M) |
| QWen-VL [5] | Openclip ViT-bigG [54] | Qwen-LM | Cross-Attention | multi-tasking datasets (Captioning, VQA, Grounding, etc. ) |
| ShareGPT4V [16] | CLIP ViT-L/14@336 | Vicuna-7B | MLP | ShareGPT4V (100K by GPT-4V, 1.2M by its learned model) |
| MobileVLM (ours) | CLIP ViT-L/14@336 | MobileLLaMA | LDP | same as LLaVA-1.5 [74] |
@@
Dataset-centric-ness. The construction of training data has become increasingly crucial. It is quite common to utilize millions of text-image pairs in the line of VLMs, where the new datasets are usually released alongside their corresponding new models. To name a few, apart from an enhanced visual receptor and novel language model called Qwen-LM [4], the multilingual multimodal Qwen-VL [5] additionally aligns the image with caption and box tuples, which sets a new record of generalist models. PALI [19] and PALI-X [18] consume an internal multi-language image-text dataset called WebLI at a scale of 12 billion. Most recently, observing the constraints of current image-text datasets like hallucination and inaccurate descriptions, ShareGPT4V [16] exploits GPT-4V [90] for generating 1.2M high-quality image-text pairs with which can surpass the LLaVA series. Similarly built with GPT-4V, LVIS-INSTRUCT4V [120] helps LLaVA-1.5 to gain substantial improvement on various VLM benchmarks.
@@ 2.5 VLM Benchmarks
Systematic and comprehensive evaluations of different VLMs are of great necessity. To this end, POPE [71] provides a benchmark for evaluating hallucination in VLMs, which formulates the evaluation as a binary classification task that prompts the VLMs to answer whether the object exists. In contrast, GQA [52] mainly centers around the VLMs’ abilities in real-world reasoning, scene understanding, and compositional question answering. TextVQA [105] consists of questions related to text in the image, evaluating the OCR and reasoning abilities of models. ScienceQA [83] consists of multimodal multiple-choice questions covering scientific topics, e.g., natural science, social science, and language science, which requires VLMs to integrate common sense into reasoning. MME [40] measures both the perception and cognition abilities of VLMs, it covers a total of 14 subtasks varying from coarse-grained to fine-grained ones. MMBench [80] is a methodically constructed multimodal dataset, which collectively covers a diverse spectrum of 20 fine-grained skills and involves a circular evaluation strategy with the incorporation of ChatGPT [88].
@@ 3.4 Efficient Projector
| Model | Common Sense Reasoning | Language Understanding | ||||||||
| ARC(c/e) | BoolQ | RTE | Winogrande | TruthfulQA | HellaSwag | PIQA | MMLU | Avg. | ||
| INCITE 3B (V1) [29] | 0.32 / 0.68 | 0.67 | 0.52 | 0.63 | 0.21 / 0.33 | 0.48 | 0.74 | 0.2675 | 0.4848 | |
| OpenLLaMA 3B (V1) [43] | 0.34 / 0.69 | 0.68 | 0.58 | 0.62 | 0.22 / 0.35 | 0.49 | 0.75 | 0.2695 | 0.4990 | |
| MobileLLaMA 2.7B | 0.32 / 0.68 | 0.61 | 0.59 | 0.63 | 0.23 / 0.36 | 0.48 | 0.75 | 0.2730 | 0.4923 | |
| TinyLLaMA 1.1B (2T) [93] | 0.23 / 0.57 | 0.59 | 0.55 | 0.57 | 0.23 / 0.39 | 0.40 | 0.70 | 0.2541 | 0.4484 | |
| Galactica 1.3B [112] | 0.28 / 0.62 | 0.62 | 0.52 | 0.55 | 0.25 / 0.41 | 0.34 | 0.63 | 0.2675 | 0.4488 | |
| OPT 1.3B [131] | 0.24 / 0.57 | 0.56 | 0.51 | 0.59 | 0.24 / 0.39 | 0.41 | 0.71 | 0.2461 | 0.4466 | |
| Pythia 1.4B [7] | 0.26 / 0.61 | 0.63 | 0.52 | 0.57 | 0.23 / 0.39 | 0.40 | 0.71 | 0.2568 | 0.4577 | |
| MobileLLaMA 1.4B | 0.26 / 0.61 | 0.53 | 0.53 | 0.59 | 0.21 / 0.35 | 0.43 | 0.73 | 0.2497 | 0.4490 | |
@@ 4.1 Training
The whole reproducible training process is composed of three stages. Firstly, we pretrain LLM foundation models on the text-only dataset RedPajama v1 [29]. Secondly, we perform supervised fine-tuning (SFT) following Vicuna [20] on a dataset of multi-turn dialogues between humans and ChatGPT from third-party platforms. Lastly, we train our vision large models using multimodal datasets.
@@
The peak learning rate is set to 3 10-4 and it decays to 3 10-5 following the cosine strategy. We warm up with 2000 iterations. We use the AdamW optimizer [82] with and and a weight decay regularization value of 0.1. The gradient clipping threshold is set to 1.0. We adopt the Pytorch lightning framework with DeepSpeed backend [99]. Specifically, we utilize ZERO 1 [98] and gradient accumulation to achieve a training speed of 18800 TGS (Tokens per second for a GPU) for the 1.4B model and 8500 TGS for the 2.7B model, on 20 nodes equipped with 8 NVIDIA Tesla A100 GPUs each.
@@ 4.2 Evaluation of MobileLLaMA
For common sense reasoning, we report the zero-shot accuracy on five prevalent benchmarks, i.e., ARCe [10], ARCc [28], BoolQ [27], Winogrande [101], and TruthfulQA [72]. Our models demonstrate strong reasoning capabilities on these benchmarks. It’s worth noting that MobileLLaMA 1.4B and 2.7B achieve the highest performance on Winogrande at both 1B and 3B levels. This indicates that our models deliver robust common sense capabilities, rather than cunningly relying on systematic biases in the datasets.
@@
| Model | Hardware | Precision |
| OpenLLaMA 3B | Snapdragon 888 | 8-bit | 3.4 | 3093 | 7.32 | 6.58 | 63.33 | ||||||||
| Snapdragon 888 | 4-bit | 2.3 | 3604 | 8.97 | 7.14 | 58.04 | |||||||||
| MobileLLaMA 2.7B | Snapdragon 888 | 8-bit | 2.7 | 3919 | 17.59 | 9.14 | 44.85 | ||||||||
| Snapdragon 888 | 4-bit | 1.5 | 3932 | 18.10 | 14.71 | 28.30 | |||||||||
| TinyLLaMA 1.1B | Snapdragon 888 | 8-bit | 1.1 | 4215 | 39.49 | 19.75 | 20.83 | ||||||||
| Snapdragon 888 | 4-bit | 0.6 | 3887 | 44.17 | 31.54 | 13.22 | |||||||||
| MobileLLaMA 1.4B | Snapdragon 888 | 8-bit | 1.4 | 3846 | 35.46 | 17.93 | 22.81 | ||||||||
| Snapdragon 888 | 4-bit | 0.7 | 3870 | 36.20 | 28.32 | 14.76 | |||||||||
| OpenLLaMA 3B | Jetson Orin | 8-bit | 3.4 | 2382 | 80.34 | 29.97 | 13.94 | ||||||||
| Jetson Orin | 4-bit | 2.3 | 3340 | 143.25 | 32.16 | 12.83 | |||||||||
| MobileLLaMA 2.7B | Jetson Orin | 8-bit | 2.7 | 3040 | 133.41 | 33.28 | 12.46 | ||||||||
| Jetson Orin | 4-bit | 1.5 | 2647 | 130.97 | 38.99 | 10.74 | |||||||||
| TinyLLaMA 1.1B | Jetson Orin | 8-bit | 1.1 | 3007 | 279.61 | 72.30 | 5.89 | ||||||||
| Jetson Orin | 4-bit | 0.6 | 3801 | 306.76 | 78.83 | 5.38 | |||||||||
| MobileLLaMA 1.4B | Jetson Orin | 8-bit | 1.4 | 3289 | 249.56 | 60.73 | 6.96 | ||||||||
@@ 4.4 MobileVLM with LoRA
| Model | LM | Hardware |
Vicuna 7B | Snapdragon | 4.70 | 8.23 | 17347 | 5.36 | 0.25 | 329.89 | ||||||||||
| OpenLLaMA 3B | Snapdragon | 2.88 | 7.98 | 27530 | 8.95 | 7.22 | 84.43 | |||||||||||
| TinyLLaMA 1B | Snapdragon | 1.18 | 7.77 | 31370 | 41.70 | 18.40 | 20.70 | |||||||||||
MobileLLaMA 2.7B | Snapdragon | 2.14 | 8.43 | 27660 | 18.36 | 12.21 | 33.10 | ||||||||||
| MobileLLaMA 1.4B | Snapdragon | 1.40 | 6.82 | 34892 | 34.93 | 21.54 | 18.51 | |||||||||||
Vicuna 7B | Jetson Orin | 4.70 | 2.89 | 9281 | 367.26 | 17.74 | 19.75 | ||||||||||
| OpenLLaMA 3B | Jetson Orin | 2.88 | 2.94 | 22270 | 474.49 | 30.66 | 12.52 | |||||||||||
| TinyLLaMA 1B | Jetson Orin | 1.18 | 2.98 | 24655 | 1253.94 | 76.63 | 5.90 | |||||||||||
MobileLLaMA 2.7B | Jetson Orin | 2.14 | 3.11 | 15678 | 440.60 | 38.34 | 8.31 | ||||||||||
@@ 4.5 Latency Measurement on Mobile Devices
| Language Model | Vision Encoder | VL Projector | Tokens | GQA | SQA | VQA | POPE | MME | MMB | ||
CLIP-B-P16-S224 [97] | MLP [75] | 196 | 55.3 | 52.8 | 38.3 | 83.2 | 1079.9 | 46.1 | ||
| CLIP-L-P14-S224 [97] | MLP [75] | 256 | 55.8 | 53.4 | 40.9 | 83.9 | 1104.4 | 50.9 | |||
| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 57.5 | 54.2 | 38.0 | 84.9 | 1123.3 | 53.9 | |||
| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 56.9 | 53.6 | 43.7 | 85.7 | 1137.7 | 52.8 | |||
| CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 | |||
CLIP-B-P16-S224 [97] | MLP [75] | 196 | 57.0 | 58.4 | 43.1 | 83.8 | 1212.2 | 54.6 | ||
| CLIP-L-P16-S224 [97] | MLP [75] | 256 | 57.9 | 58.6 | 45.3 | 85.1 | 1285.0 | 57.7 | |||
| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 59.5 | 58.9 | 43.9 | 85.1 | 1275.6 | 59.5 | |||
| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 59.1 | 58.3 | 47.3 | 85.8 | 1333.1 | 57.1 | |||

### v2

@@ 2.3 VLMs
| Model | Vision Encoder | Language Model |
Cross-modality Design |
Multimodal Training Corpora |
| CLIP [97] | ViT, ResNet | Transformer |
Linear-Projection |
WebImageText [97] (400M image-text pairs) |
| BLIP [67] | ViT | MED |
Cross-Attention |
| BLIP-2 [66] | CLIP/Eva-CLIP ViT | OPT, Flan-T5 |
Q-Former |
same as BLIP |
| InstructBLIP [30] | ViT-G/14 | Flan-T5, Vicuna |
Q-Former w/ FC |
13 held-in out of 26 public datasets |
| Flamingo [2] | NFNet-F6 [59] | Chinchilla |
Perceiver-Resampler [57] |
M3W(43M), ALIGN(1.4B) [58], LTIP (312M), VTP (27M) |
| LLaVA [76] | CLIP ViT-L/14 | Vicuna 7B/13B |
Linear-Projection |
filtered CC-595K from CC3M [103], LLaVA-Instruct-158K |
| LLaVA-1.5 [74] | CLIP ViT-L@336 | Vicuna-7B/13B |
MLP |
a subset of InstructBLIP (1.2M) |
| MiniGPT-4 [133] | EVA-CLIP ViT-G/14 | Vicuna-7B |
Q-Former |
LAION, CC, SBU [91] |
| Shikra [15] | CLIP ViT-L/14 | Vicuna-7/13B |
FC-layer |
| mPLUG-Owl [126] | CLIP ViT-L/14 | LLaMA-7B |
Cross-Attention |
LAION-400M, COYO-700M [11], CC, COCO |
| KOSMOS-2 [94] | VLMo [6] | MAGNETO [119] |
Perceiver-Resampler |
GRIT (curated with LAION-2B, COYO-700M) |
| QWen-VL [5] | Openclip ViT-bigG [54] | Qwen-LM |
Cross-Attention |
multi-tasking datasets (Captioning, VQA, Grounding, etc. ) |
| ShareGPT4V [16] | CLIP ViT-L/14@336 | Vicuna-7B |
MLP |
ShareGPT4V (100K by GPT-4V, 1.2M by its learned model) |
| MobileVLM (ours) | CLIP ViT-L/14@336 | MobileLLaMA |
LDP |
same as LLaVA-1.5 [74] |
@@
Dataset-centric-ness. The construction of training data has become increasingly crucial. It is common to utilize millions of text-image pairs in the line of VLMs, where the new datasets are usually released alongside their corresponding new models. To name a few, apart from an enhanced visual receptor and novel language model called Qwen-LM [4], the multilingual multimodal Qwen-VL [5] additionally aligns the image with caption and box tuples, which sets a new record of generalist models. PALI [19] and PALI-X [18] consume an internal multi-language image-text dataset called WebLI at a scale of 12 billion. Most recently, observing the constraints of current image-text datasets like hallucination and inaccurate descriptions, ShareGPT4V [16] exploits GPT-4V [90] for generating 1.2M high-quality image-text pairs with which can surpass the LLaVA series. Similarly built with GPT-4V, LVIS-INSTRUCT4V [120] helps LLaVA-1.5 to gain substantial improvement on various VLM benchmarks.
@@ 2.5 VLM Benchmarks
Systematic and comprehensive evaluations of different VLMs are of great necessity. To this end, POPE [71] provides a benchmark for evaluating hallucination in VLMs, which formulates the evaluation as a binary classification task that prompts the VLMs to answer whether the object exists. In contrast, GQA [52] mainly centers around the VLMs’ abilities in real-world reasoning, scene understanding, and compositional question answering. TextVQA [105] consists of questions related to the text in the image, evaluating the OCR and reasoning abilities of models. ScienceQA [83] consists of multimodal multiple-choice questions covering scientific topics, e.g., natural science, social science, and language science, which requires VLMs to integrate common sense into reasoning. MME [40] measures both the perception and cognition abilities of VLMs, it covers a total of 14 subtasks varying from coarse-grained to fine-grained ones. MMBench [80] is a methodically constructed multimodal dataset, which collectively covers a diverse spectrum of 20 fine-grained skills and involves a circular evaluation strategy with the incorporation of ChatGPT [88].
@@ 3.4 Efficient Projector
| Model | Common Sense Reasoning | Language Understanding | |||||||
| ARC | BoolQ | RTE | Winogrande | TruthfulQA | HellaSwag | PIQA | MMLU | Avg. | |
| INCITE 3B (V1) [29] | 0.32 / 0.68 | 0.67 | 0.52 | 0.63 | 0.21 / 0.33 | 0.48 | 0.74 | 0.2675 | 0.4848 |
| OpenLLaMA 3B (V1) [43] | 0.34 / 0.69 | 0.68 | 0.58 | 0.62 | 0.22 / 0.35 | 0.49 | 0.75 | 0.2695 | 0.4990 |
| MobileLLaMA 2.7B | 0.32 / 0.68 | 0.61 | 0.59 | 0.63 | 0.23 / 0.36 | 0.48 | 0.75 | 0.2730 | 0.4923 |
| TinyLLaMA 1.1B (2T) [93] | 0.23 / 0.57 | 0.59 | 0.55 | 0.57 | 0.23 / 0.39 | 0.40 | 0.70 | 0.2541 | 0.4484 |
| Galactica 1.3B [112] | 0.28 / 0.62 | 0.62 | 0.52 | 0.55 | 0.25 / 0.41 | 0.34 | 0.63 | 0.2675 | 0.4488 |
| OPT 1.3B [131] | 0.24 / 0.57 | 0.56 | 0.51 | 0.59 | 0.24 / 0.39 | 0.41 | 0.71 | 0.2461 | 0.4466 |
| Pythia 1.4B [7] | 0.26 / 0.61 | 0.63 | 0.52 | 0.57 | 0.23 / 0.39 | 0.40 | 0.71 | 0.2568 | 0.4577 |
| MobileLLaMA 1.4B | 0.26 / 0.61 | 0.53 | 0.53 | 0.59 | 0.21 / 0.35 | 0.43 | 0.73 | 0.2497 | 0.4490 |
@@ 4.1 Training
The whole reproducible training process is composed of three stages. Firstly, we pre-train LLM foundation models on the text-only dataset RedPajama v1 [29]. Secondly, we perform supervised fine-tuning (SFT) following Vicuna [20] on a dataset of multi-turn dialogues between humans and ChatGPT from third-party platforms. Lastly, we train our vision large models using multimodal datasets.
@@
The peak learning rate is set to 3 10 and it decays to 3 10 following the cosine strategy. We warm up with 2000 iterations. We use the AdamW optimizer [82] with and and a weight decay regularization value of 0.1. The gradient clipping threshold is set to 1.0. We adopt the Pytorch lightning framework with DeepSpeed backend [99]. Specifically, we utilize ZERO 1 [98] and gradient accumulation to achieve a training speed of 18800 TGS (Tokens per second for a GPU) for the 1.4B model and 8500 TGS for the 2.7B model, on 20 nodes equipped with 8 NVIDIA Tesla A100 GPUs each.
@@ 4.2 Evaluation of MobileLLaMA
For common sense reasoning, we report the zero-shot accuracy on five prevalent benchmarks, i.e., ARC [10], ARC [28], BoolQ [27], Winogrande [101], and TruthfulQA [72]. Our models demonstrate strong reasoning capabilities on these benchmarks. It’s worth noting that MobileLLaMA 1.4B and 2.7B achieve the highest performance on Winogrande at both 1B and 3B levels. This indicates that our models deliver robust common sense capabilities, rather than cunningly relying on systematic biases in the datasets.
@@
| Model | Hardware | Precision | Size (GB) | (tokens/s) | (tokens/s) | (tokens/s) | (s) |
| OpenLLaMA 3B | Snapdragon 888 | 8-bit | 3.4 | 3093 | 7.32 | 6.58 | 63.33 |
| Snapdragon 888 | 4-bit | 2.3 | 3604 | 8.97 | 7.14 | 58.04 | |
| MobileLLaMA 2.7B | Snapdragon 888 | 8-bit | 2.7 | 3919 | 17.59 | 9.14 | 44.85 |
| Snapdragon 888 | 4-bit | 1.5 | 3932 | 18.10 | 14.71 | 28.30 | |
| TinyLLaMA 1.1B | Snapdragon 888 | 8-bit | 1.1 | 4215 | 39.49 | 19.75 | 20.83 |
| Snapdragon 888 | 4-bit | 0.6 | 3887 | 44.17 | 31.54 | 13.22 | |
| MobileLLaMA 1.4B | Snapdragon 888 | 8-bit | 1.4 | 3846 | 35.46 | 17.93 | 22.81 |
| Snapdragon 888 | 4-bit | 0.7 | 3870 | 36.20 | 28.32 | 14.76 | |
| OpenLLaMA 3B | Jetson Orin | 8-bit | 3.4 | 2382 | 80.34 | 29.97 | 13.94 |
| Jetson Orin | 4-bit | 2.3 | 3340 | 143.25 | 32.16 | 12.83 | |
| MobileLLaMA 2.7B | Jetson Orin | 8-bit | 2.7 | 3040 | 133.41 | 33.28 | 12.46 |
| Jetson Orin | 4-bit | 1.5 | 2647 | 130.97 | 38.99 | 10.74 | |
| TinyLLaMA 1.1B | Jetson Orin | 8-bit | 1.1 | 3007 | 279.61 | 72.30 | 5.89 |
| Jetson Orin | 4-bit | 0.6 | 3801 | 306.76 | 78.83 | 5.38 | |
| MobileLLaMA 1.4B | Jetson Orin | 8-bit | 1.4 | 3289 | 249.56 | 60.73 | 6.96 |
@@ 4.4 MobileVLM with LoRA
| Model | LM | Hardware | Size (GB) | (ms/patch) | (tokens/s) | (tokens/s) | (tokens/s) | (s) |
| LLaVA-v1.5-336 | Vicuna 7B | Snapdragon | 4.70 | 8.23 | 17347 | 5.36 | 0.25 | 329.89 |
| OpenLLaMA 3B | Snapdragon | 2.88 | 7.98 | 27530 | 8.95 | 7.22 | 84.43 | |
| TinyLLaMA 1B | Snapdragon | 1.18 | 7.77 | 31370 | 41.70 | 18.40 | 20.70 | |
| MobileVLM-336 | MobileLLaMA 2.7B | Snapdragon | 2.14 | 8.43 | 27660 | 18.36 | 12.21 | 33.10 |
| MobileLLaMA 1.4B | Snapdragon | 1.40 | 6.82 | 34892 | 34.93 | 21.54 | 18.51 | |
| LLaVA-v1.5-336 | Vicuna 7B | Jetson Orin | 4.70 | 2.89 | 9281 | 367.26 | 17.74 | 19.75 |
| OpenLLaMA 3B | Jetson Orin | 2.88 | 2.94 | 22270 | 474.49 | 30.66 | 12.52 | |
| TinyLLaMA 1B | Jetson Orin | 1.18 | 2.98 | 24655 | 1253.94 | 76.63 | 5.90 | |
| MobileVLM-336 | MobileLLaMA 2.7B | Jetson Orin | 2.14 | 3.11 | 15678 | 440.60 | 38.34 | 8.31 |
@@ 4.5 Latency Measurement on Mobile Devices
| Language Model | Vision Encoder | VL Projector | Tokens | GQA | SQA | VQA | POPE | MME | MMB |
| MobileLLaMA 1.4B | CLIP-B-P16-S224 [97] | MLP [75] | 196 | 55.3 | 52.8 | 38.3 | 83.2 | 1079.9 | 46.1 |
| CLIP-L-P14-S224 [97] | MLP [75] | 256 | 55.8 | 53.4 | 40.9 | 83.9 | 1104.4 | 50.9 | |
| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 57.5 | 54.2 | 38.0 | 84.9 | 1123.3 | 53.9 | |
| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 56.9 | 53.6 | 43.7 | 85.7 | 1137.7 | 52.8 | |
| CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 | |
| MobileLLaMA 2.7B | CLIP-B-P16-S224 [97] | MLP [75] | 196 | 57.0 | 58.4 | 43.1 | 83.8 | 1212.2 | 54.6 |
| CLIP-L-P16-S224 [97] | MLP [75] | 256 | 57.9 | 58.6 | 45.3 | 85.1 | 1285.0 | 57.7 | |
| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 59.5 | 58.9 | 43.9 | 85.1 | 1275.6 | 59.5 | |
| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 59.1 | 58.3 | 47.3 | 85.8 | 1333.1 | 57.1 | |